OpenStudy (anonymous):

What are the real or imaginary solutions of the polynomials x^4 -52x^2 + 576 = 0 A. No solution B.-4,-6 D.-4,4,-6,6

OpenStudy (anonymous):

\[x^4-52x^2+576=0\] A. No solution B.-4,-6 D.-4,4,-6,6

OpenStudy (anonymous):

Ok I got it, it's D (x+4)(x−4)(x+6)(x−6)=0 x+4=0 or x−4=0 or x+6=0 or x−6=0 x=−4 or x=4 or x=−6 or x=6 Answer: x=−4 or x=4 or x=−6 or x=6